OpenStudy (anonymous):

452.7 x 10^3 - 2.7 x 10^5 in scientific notation

jagr2713 (jagr2713):

Do you know anything about scientific notation?

OpenStudy (paxpolaris):

457x10^3 =4.527.x10^5

OpenStudy (paxpolaris):

\[\ \ \ \ 4.527\times 10^5 \\ \underline{-2.700\times10^5}\]

OpenStudy (paxpolaris):

\[\ \ \ \ 4.527\times 10^5 \\ \underline{-2.700\times10^5} \\ \underline{\ \ \ \color{green}{1.827\times 10^5}}\]